OpenStudy (anonymous):

Solve the problem. (Must show work!) 1. A job pays $56 for 8 hours of work. How much money does the job pay per hour?

OpenStudy (anonymous):

7

OpenStudy (texaschic101):

its another proportion connor... 56/8 = x/1 cross multiply (8)(x) = (56)(1) 8x = 56 divide both sides by 8 x = 7

OpenStudy (anonymous):

@texaschic101 Thanks, but I accidentally pose this question.